12) Using Booths Algorithm, solve (-16) X(+5). Fill in the details in the below table. (40 points) Count A Q Q M Comments
Answer #1

Q А count M Comments 110000 M O 6 000 lot we addi 001000 000010 110000 Q.Q-, 210 so we subtract and then sight shift and decrAfter the last right shift will get Q As the count becomes 111110 11 0000 AQ is the answer le this is as compliment 111110 1We will have to iterate till count times count is the number of bits in the multiplicand.
Here we take 6 bits as we have to do 2's compliment of 16.
There will be 3 conditions :
1) if Q0Q-1 = 01 then we will subtract M from A then we right shift and decrement the count.
2) if Q0Q-1 = 10 then we will add the M to A then we right shift and decrement the count.
3) if Q0Q-1 = 00 or 11 then we will just right shift and decrement the count.​​​​​​

PLEASE DON'T FORGET TO LIKE...